3. For each of the following distributions, find the expected value of X (a) f(z,8) = θ(1-9)-, z = 1, 2, . . . . 0 (b) f(z,8) = (θ + 1)2 , z > 1,0 > 0 (e) f(z,8) = θ2ze-e, z > 0, θ > 0 9s 1 -0-2
Answer #1

a)

E(X)=\sum_{x=0}^{\infty }x*\theta * (1-\theta )^{x-1}

\Rightarrow E(X)=\theta \sum_{x=0}^{\infty }x* (1-\theta )^{x-1}

\Rightarrow E(X)=\theta [1+2*(1-\theta )+3*(1-\theta )^2+...]

\Rightarrow E(X)=\theta *(1-(1-\theta ))^{-2}

\Rightarrow E(X)=\theta *\frac{1}{\theta ^2}=\frac{1}{\theta }

\Rightarrow E(X)=\frac{1}{\theta }

b) E(X)=\int_{1}^{\infty }x*(\theta+1) * x^{-\theta -2}dx

\Rightarrow E(X)=(\theta+1) \int_{1}^{\infty }x^{-\theta -1}dx

\Rightarrow E(X)=(\theta+1)*[\frac{x^{-\theta }}{-\theta }]_{1}^{\infty }

\Rightarrow E(X)=(\theta+1)*[\frac{1}{\theta }]

\Rightarrow E(X)=1+\frac{1}{\theta }

c) E(X)=\int_{0}^{\infty }x*\theta ^2*xe^{-\theta x}dx

\Rightarrow E(X)=\theta ^2\int_{0}^{\infty }x^2*e^{-\theta x}dx

\Rightarrow E(X)=\theta ^2*\frac{\Gamma (3)}{\theta^3}=\frac{2}{\theta }

\Rightarrow E(X)=\frac{2}{\theta }